how many seconds would it a take a ball to fall 225 feet?
Answer by stanbon(75887)

how many seconds would it a take a ball to fall 225 feet?
----
16t^2 = 225
t^2 = 225/16
t = 15/4 = 3 3/4 seconds
============================
Cheers,
Stan H.